The Compound interest is 1590 for 2 years at the rate of 12% PA Find the simple interest for the same period for 16%
Hint Answer is 2000
Compound interest=1590 it is not the principal
Rate =12%
Time=2years

Answer by stanbon(75887)

The Compound interest is 1590 for 2 years at the rate of 12% PA Find the simple interest for the same period for 16%
Hint Answer is 2000
Compound interest=1590 it is not the principal
Rate =12%
Time=2years
----------
Compound information:
P+I = P(1+(r/n))^(nt)
P+1590 = P(1+(0.12/1))^(1*2)
P+1590 = P(1.2544)
0.2544P = 1590
P = $6250
-----
Simple Interest information:
I = Prt
I = 6250*0.16*2
---
I = $2000
=============
Cheers,
Stan H.
=============